OpenStudy (anonymous):

find all zeros of P, real and complex. Factor P completely. P(x)= x^4 +6x^2 +9

OpenStudy (akashdeepdeb):

\[P(x)= x^4 +6x^2 +9\] Now P(x) is a bi-quadratic polynomial BUT we CAN write it as a quadratic polynomial by taking x^2 = y Now \[ P(y) = y^2 + 6y + 9 \] Where y = x^2 Now find out the roots of P(y) = -3,-3 y = x^2 = -3 x = +√3i , -√3i y = x^2 = -3 x = +√3i , - √3i So the roots pf P(x) are --> +√3i , -√3i, +√3i , - √3i Understood? :)

OpenStudy (anonymous):

why do we find the roots of -3?

OpenStudy (akashdeepdeb):

Because ORIGINALLY we need to find the roots of x and NOT y.

OpenStudy (anonymous):

ooohh okay thank you so much! I don't know if clicking the best response helps you in any way but I clicked it just in case!

OpenStudy (akashdeepdeb):

Helps a lot. You're welcome. :)